Need this to study for a test on Friday. Thank you!
A Quadratic relation had roots 0 and -6 and a maximum at (-3, 4). Determine the equation of the relation.

Answer by ewatrrr(24785)

V(-3,4) max
y = a(x+3)^2 + 4
0 = a(3)^2 + 4
-4/9 = a
.......
0 = a(-3)^2 + 4
-4/9 = a
........
y = (-4/9)(x+3)^2 + 4